2. A new non-invasive screening test is proposed that is claimed to be able to identify patients with The new to uestions they are sample impaired glucose tolerance base each patients responses to the submits a blood on patient also below. (10 points) test given to impa red tolerance. Each tabulated tive tive for determined. The results are Impai nega status is Not and their glucose tolerance Tolerance 13 Screening Test impaired Glucose Negative a) What is the sensitivity of the screening test? b) What is the false positive fraction of the screening test?
Answer #1

In this case –

Screening Test

Impaired Glucose Tolerance

Not Impaired

Positive

17 (True positive)

13 (False Positive)

Negetive

8 (False Negetive)

37 (True Negetive)

a) Sensitivity of the screening test –

TP/(TP+FN) = 17/25 = 0.68

b) False positive fraction of screening test –

1- specificity = 1 – [TN/(TN+FP)] = 1 – [37/50] = 1- 0.74 = 0.26